# Uniform approximation of Betti numbers

### v1

@@ 2. Perverse sheaves
As a consequence, Tate twists $M\mapsto M(n)$ are $t$-exact functors. Any constructible object $F$ of $D(X)$ is bounded. Indeed, if we define $f$ as the coproduct of the embeddings of the strata over which $F$ becomes dualizable, then $f^{*}(F)$ is dualizable in $D(X^{\prime})$, and since $f^{*}$ is conservative (by localization) and $t$-exact, this proves the claim. Remark that condition (f) is only reasonnable with $\mathbf{F}$-linear coefficients for some field $\mathbf{F}$. [^3] This restriction comes from Condition (f). It can be avoided but we would need two $t$-structures whose restrictions to dualizable objects would be exchanged through duality and that would only make notations worse. We denote by $D^{\leq i}_{c}(X)$ (by $D^{\geq i}_{c}(X)$ respectively) the full subcategory of $D_{c}(X)$ that consists of constructible objects with cohomology concentrated in degrees $\leq i$ (in degrees $\geq i$, respectively) with respect to the $t$-structure above. Such a $t$-structure will be called ordinary.
@@
It is well known that, to say the least, the existence of such a $t$-structure is very conjectural as far as $\mathit{DM}$ is concerned. However, $D(X)=D(X,\mathbf{Q}_{\ell})$ and $D(X)=D(X,\mathbf{F}_{\mathcal{U}})$ have canonical $t$-structures coming from the usual $t$-structures on the derived category of sheaves of $\Lambda$-modules on the small étale site of $X$, with $\Lambda$ and finite commutative ring. If moreover $X$ is noetherian, all the properties listed above will be fulfilled in these cases.
@@
    For any affine immersion $i\colon U\to X$, the functor $i_{!}\colon D(U)\to D(X)$ is $t$-exact with respect to the perverse $t$-structure;
@@
Left $F$ be in $^{p}D^{\leq 0}(X)$. We choose $k\colon S\to X$ a nowhere dense closed immersion with open complement $U$ whose reduction is regular, so that the restriction of $F$ on $U$ is dualizable and in $D^{\leq\dim U}(U)$ (here the dimesion is a locally constant function on $U$). Then the restriction of $F$ on $S$ is in $^{p}D^{\leq 0}(S)$ by definition. If $Z$ is nowhere dense in $X$, we may choose $U$ away from $Z$. This means that $Z$ belongs to $S$ with $\dim S<\dim X$, and we conclude by induction on the dimension. Otherwise, if $d=\dim(U\cap Z)$, the restriction of $F$ on $Z\cap S$ is in $^{p}D^{\leq 0}(S\cap Z)$ (by induction as above) and the restriction of F[d]F[-d] on $Z\cap U$ is dualizable and in $D^{\leq 0}(Z\cap U)$. Therefore, the restriction of $F$ on $Z$ is in $^{p}D^{\leq 0}(Z)$. ∎
@@
$D(-,\mathbf{F})$ with $\mathbf{F}$ a finite field, $D(-,\mathbf{Q}_{\ell})$ with $\ell$ any prime number, and $D(-,\mathbf{F}_{\mathcal{U}})$ with $\mathcal{U}$ any ultrafilter on the set of prime numbers, as defined in 1.3 and 1.4 are Artin-Verdier six-functors formalisms. Other examples are the de Rham six-functors formalism, the $\mathbf{F}$-linear version of Betti six-functors formalism (for any field of coefficient $\mathbf{F}$), and the rigid six-functors formalism.
@@
    For any quasi-finite morphism $f\colon X\to Y$, the functor $f^{*}\colon D(_{c}Y)\to D_{c}(X)$ is right $t$-exact and the exceptional pullback functor $f^{!}\colon D_{c}(Y)\to D_{c}(X)$ is left $t$-exact.
@@
    For any quasi-finite morphism $f\colon X\to Y$, the functor $f_{!}\colon D(_{c}Y)\to D_{c}(X)$ is right $t$-exact and $f_{*}\colon D_{c}(Y)\to D_{c}(X)$ is left $t$-exact.
@@
To prove (a), we use criterium (ii) of Proposition 2.15. Let $T$ be a closed subscheme of $X$. There is a dense open subscheme $U$ of the closure of the image of $T$ in $Y$ such that $\imath_{U}^{*}(F)$ belongs to $D^{\leq\dim U}_{c}(U)$. Therefore, if $S$ denotes the pullback of $U$ in $T$, we have $\imath_{U}^{*}\,f^{*}(F)$ in $D^{\leq\dim U}_{c}(S)$. Since $\dim S=\dim U$ (because $f$ is quasi-finite), this proves that $f^{*}(F)$ is in $^{p}D^{\leq 0}(X)$. By Verdier duality, this proves that $f^{!}$ is left $t$-exact. Property (b) follows straight away from property (a). ∎
@@ 3. Uniformly perverse motivic generators
$\mathit{rk}(X,Z)=(-1)^{d}\mathit{Tr}(1_{M(X,Z)})\in\mathbf{Z}$
@@ A Lefchetz filtration of $X$ is a finite sequence of closed subschemes of the form
(where the cohomology is defined using the ordinary $t$-structure as in Remark 2.21). The rank of $H^{d}(X,j_{!}\mathbf{1}_{X-Z})$ (defined as the trace of its identity) is equal to the class of $\mathit{rk}(X,Z)$ in $\pi_{0}\operatorname{Map}_{D(k)}(\mathbf{1},\mathbf{1})$.
@@
that commute with the six operations. If $M(X,Z)^{\wedge}$ denotes the dual of $M(X,Z)$ in $\mathit{DM}(\mathrm{Spec}(k))$, we thus have (with the notations of Remark 2.21)
@@
$\mathit{DM}_{c}(Y)\to D^{b}_{c}(Y,\mathbf{F}_{\ell})\quad\text{and}\quad\mathit{DM}_{c}(Y)\to D^{b}_{c}(Y,\mathbf{Q}_{\ell})$
@@
respectively. Therefore, by virtue of Theorem 2.22 for both $D(-,\mathbf{Q}_{\ell})$ and $D(-,\mathbf{F}_{\ell})$, we see that we can apply the discussion above with $C$ be the $\ell$-adic realization of the motive $a_{*}j_{!}j^{*}\mathbf{Z}_{X}[d]$ in the stable $\infty$-category $D^{b}_{c}(\mathrm{Spec}(k),\mathbf{Z}_{\ell})\cong\mathit{Perf}(\mathbf{Z}_{\ell})$. This proves the last part of the lemma. ∎

### v2

@@ 2. Perverse sheaves
As a consequence, Tate twists $M\mapsto M(n)$ are $t$-exact functors. Any constructible object $F$ of $D(X)$ is bounded. Indeed, if we define $f$ as the coproduct of the embeddings of the strata over which $F$ becomes dualizable, then $f^{*}(F)$ is dualizable in $D(X^{\prime})$, and since $f^{*}$ is conservative (by localization) and $t$-exact, this proves the claim. Remark that condition (f) is only reasonnable with $\mathbf{F}$-linear coefficients for some field $\mathbf{F}$. [^4] Condition (f) can be avoided but we would need two $t$-structures whose restrictions to dualizable objects would be exchanged through ordinary duality, and that would only make notations worse. We denote by $D^{\leq i}_{c}(X)$ (by $D^{\geq i}_{c}(X)$ respectively) the full subcategory of $D_{c}(X)$ that consists of constructible objects with cohomology concentrated in degrees $\leq i$ (in degrees $\geq i$, respectively) with respect to the $t$-structure above. Such a $t$-structure will be called ordinary.
@@
It is well known that, to say the least, the existence of such a $t$-structure is very conjectural as far as $\mathit{DM}$ is concerned. However, $D(X)=D(X,\mathbf{Q}_{\ell})$, $D(X)=D(X,\mathbf{F}_{\ell})$ and $D(X)=D(X,\mathbf{F}_{\mathcal{U}})$ have canonical $t$-structures coming from the usual $t$-structures on the derived category of sheaves of $\Lambda$-modules on the small étale site of $X$, with $\Lambda$ and finite commutative ring. If moreover $X$ is noetherian, all the properties listed above will be fulfilled in these cases.
@@
    For any closed immersion $i\colon Z\to X$, the functor $i_{!}\colon D(Z)\to D(X)$ is $t$-exact with respect to the perverse $t$-structure, and for any open immersion $j\colon U\to X$, the functor $j^{*}\colon D(X)\to D(U)$ is $t$-exact with respect to the perverse $t$-structure;
@@
This is obvious since its right adjoint $i_{*}$ is (left) $t$-exact with respect to the perverse $t$-structure. ∎
@@
$D(-,\mathbf{F})$ with $\mathbf{F}$ a finite field, $D(-,\mathbf{Q}_{\ell})$ with $\ell$ any prime number, and $D(-,\mathbf{F}_{\mathcal{U}})$ with $\mathcal{U}$ any ultrafilter on the set of prime numbers, as defined in 1.3 and 1.4 are Artin-Verdier six-functors formalisms – properties (i) and (ii) are even known to hold over any quasi-excellent base, after the work of Gabber [ILO14, Exposé XVII, Théorème 0.2 and Exposé XV, Théorème 1.1.2]. Other examples are the de Rham six-functors formalism, the $\mathbf{F}$-linear version of Betti six-functors formalism (for any field of coefficient $\mathbf{F}$), and the rigid six-functors formalism.
@@
    For any quasi-finite morphism $f\colon X\to Y$, the functor $f^{*}\colon D_{c}(Y)\to D_{c}(X)$ is right $t$-exact and the exceptional pullback functor $f^{!}\colon D_{c}(Y)\to D_{c}(X)$ is left $t$-exact.
@@
    For any quasi-finite morphism $f\colon X\to Y$, the functor $f_{!}\colon D_{c}(Y)\to D_{c}(X)$ is right $t$-exact and $f_{*}\colon D_{c}(Y)\to D_{c}(X)$ is left $t$-exact.
@@
To prove (a), we use criterium (ii) of Proposition 2.15. Let $T$ be a closed subscheme of $X$. There is a dense open subscheme $U$ of the closure of the image of $T$ in $Y$ such that $\imath_{U}^{*}(F)$ belongs to $D^{\leq\dim U}_{c}(U)$. Therefore, if $S$ denotes the pullback of $U$ in $T$, we have $\imath_{U}^{*}\,f^{*}(F)$ in $D^{\leq\dim U}_{c}(S)$. Since $\dim S=\dim U$ (because $f$ is quasi-finite and $S$ dominates $U$ by construction), this proves that $f^{*}(F)$ is in $^{p}D^{\leq 0}(X)$. By Verdier duality, this proves that $f^{!}$ is left $t$-exact. Property (b) follows straight away from property (a). ∎
@@ 3. Uniformly perverse motivic generators
$\mathit{rk}(X,Z)=(-1)^{d}\mathit{Tr}\big(1_{M(X,Z)}\big)\in\mathbf{Z}$
@@ A Lefchetz filtration of $X$ is a finite sequence of closed subschemes of the form
(where the cohomology is defined using the ordinary $t$-structure as in Remark 2.21). The rank of $H^{d}(X,j_{!}\mathbf{1}_{X-Z})$ (defined as the trace of its identity) is equal to the class of $\mathit{rk}(X,Z)$ in $\pi_{0}\operatorname{Map}_{D(\mathrm{Spec}(k))}(\mathbf{1},\mathbf{1})$.
@@
that commute with the six operations. With the notations of Remark 2.21, if $M(X,Z)^{\wedge}\cong a_{*}j_{!}j^{*}\mathbf{Z}_{X}$ denotes the dual of $M(X,Z)$ in $\mathit{DM}(\mathrm{Spec}(k))$, we thus have
@@
$\mathit{DM}_{c}(Y)\to D^{b}_{c}(Y,\mathbf{F}_{\ell})\quad\text{and}\quad\mathit{DM}_{c}(Y)\to D^{b}_{c}(Y,\mathbf{Z}_{\ell})\otimes\mathbf{Q}=D^{b}_{c}(Y,\mathbf{Q}_{\ell})$
@@
respectively. Therefore, by virtue of Theorem 2.22 for $D^{b}_{c}(-,\mathbf{Q}_{\ell})$ and $D^{b}_{c}(-,\mathbf{F}_{\ell})$, we see that we can apply the discussion above with $C$ be the $\ell$-adic realization of the motive $a_{*}j_{!}j^{*}\mathbf{Z}_{X}[d]$ in the stable $\infty$-category $D^{b}_{c}(\mathrm{Spec}(k),\mathbf{Z}_{\ell})\cong\mathit{Perf}(\mathbf{Z}_{\ell})$. This proves the last part of the lemma. ∎